43. In a region of chromosome 4 there are three genes, j, k, 1. 10 cM 20 cM The coefficient of coincidence in this region is 0.8, and it is known that the interference is driven by j-k recombination where recombination between j and k interferes with recombination in the region between k and 1. A true -breeding k is mated with a true-breeding j 1 individual. The resultant triple heterozygote is test crossed. What classes (include phenotypes and proportions) are expected in the next generation if 1000 offspring?
Answer #1

The steps involved in this problem:

1. Calculation of DCO from COC
2. Calculation of SCO and
3. Calculation of PCO.

1. Calculation of DCO freequency :
COC = 0.8.
COC = Prop of ODCO/ Prop of EDCO.
=> 0.8 = Prof of ODCO/ (10/100 X 20/100) => DCO = 0.8 X 0.1 X 0.2 = 0.016
0.016

The Freequency of DCO = 0.016 *1000 = 16.


2. Calculation of SCO:.
The Distiance between two genes= Prop of SCO+ Prop of DCO.
So, Prop of SCO = Distiance between two genes - Prop of DCO.
The freequency = Prop of SCO * Total no of progeny


Between first pair of genes =
The Distance = 10, The proportion = 10/100 = 0.1
Prop of SCO = 0.1- 0.016 = 0.084
The freequency = 0.084 * 1000 = 84.


Between second pair of genes =
The Distance = 20, The proportion = 20/100 = 0.20
Prop of SCO = 0.20- 0.016 = 0.184
The freequency = 0.184 * 1000 = 184.

3. Calculation of PCO:
Parental combinations = Total freequency - SCO - DCO
=> 1000 - 184 - 84- 16 = 716